The length of the first square is 3 times more than the length of the second square. The total area is 149 in2.: The length of the first square is 3 times more than the length of the second square. The total area is 149 in2.
Answer by checkley77(3654)

x^2+(3x)^2=149
x^2+9x^2=149
10x^2=149
x^2=149/10
x^2=14.9
x=3.86 in is the side of the smaller square
3*3.86=11.58 in for the side of the larger square.
Proof:
3.86^2+11.58^2=149
14.9+134.1=149
149=149